The measure of an angle increased by 20 degrees is equal to the measure of its compliment
Answer by macston(5194)

A=angle; C=complement=A+20;
.
A+C=90
A+(A+20)=90
2A=70
A=35
.
ANSWER: The angle is 35 degrees.
.
CHECK:
.
C=A+20=35+20=55
Complement=55 degrees
.
A+C=90
36+66=90
90=90
.
NOTE: Complement (as in complementary angles) is not
spelled the same as compliment (polite expression of praise).